A soap film of refractive index 1.43is illuminated by white light incident at an angle 300. The refracted light is examined by a spectroscope in which dark band corresponding to wavelength 6х10-7m is observed. Calculate the thickness of the film.


1
Expert's answer
2021-04-15T07:16:48-0400

"μ = 1.43 \\

λ = 6 \times 10^{-7} \;m \\

i = 30º"

Since the dark band was found n=1 condition for darkness in reflected light is given by

"2μ \times t \times cosγ = nλ \\

μ = \frac{sini}{cosγ} \\

sinγ = \frac{sini}{μ} \\

= \frac{sin30}{1.43} \\

= 0.349 \\

sin^2γ = 0.1222 \\

1 -cos^2γ = 0.1222 \\

cos^2γ = 0.8788 \\

cosγ = 0.93744"

Thickness is found by the equation:

"t = \frac{nλ}{2μcosγ} \\

= \frac{1 \times 6 \times 10^{-7}}{2 \times 1.43 \times 0.93744} \\

= 2.23 \times 10^{-7} \;m"